OpenStudy (anonymous):

Evaluate u + xy, for u = 20, x = 9, and y = 8.

OpenStudy (ajprincess):

jst plug in the values. u + xy=20+(9*8)=?

OpenStudy (anonymous):

20 + 9*8

OpenStudy (anonymous):

=92

OpenStudy (anonymous):

Thanks a lot!